There is a problem that asks: what is the value of y?

It is a hexagon with an angle that are 130 degrees, 75, 145, 115, 125. Then there is a exterior angle stick out of the side and it is y. I need the value of y.

Answer by Edwin McCravy(20055)

I'm guessing your hexagon looks something like this:

 

The sum of the interior angles of a polygon is (n-2)·180°.

A hexagon has 6 sides so n=6 and the sum of the interior angles is 

(6-2)·180° = 4·180° = 720°

The sum of the 5 given angles is 130°+75°+145°+115°+125° = 590°

So the remaing 6th interior angle is 720°-590° = 130°,
so we will label it 130° in red. 



And now y° and 130° are supplementary and have sum 180°.

Therefore y° = 180°-130° = 50° 

Edwin